Fermi National Laboratory
 

Operation Summary for February 22, 2010 - February 24, 2010

 

Monday Feb. 22

The day shift began with Operations monitoring the Tevatron (TeV) store 7628 (39.56E30) and the Recycler stash (367.8E10), with the Antiproton Source stacking (11.8E10), with the Linac, Booster, and Main Injector operating normally, with the NuMI, MiniBooNE, and Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beam.

During the day shift, Operations terminated store 7628 at 10 AM and put the TeV into HEP shot setup. Operations established store 7629 at 11:30 AM with an Average Initial Luminosity (AIL) of 248.36E30. The low initial luminosity was due to an emittance blow-up and beam position problems in the Recycler. MTA personnel conducted beam studies during the morning. A Booster RF station (BRF8) caused problems through the day shift.

During the evening shift, Operations disabled the MTest beam at 5:40 PM. Operators turned off BRF8 due to modulator issues.

Tuesday Feb. 23

The midnight shift began with Operations monitoring store 7629 (251.57E30) and the stash (127.6E10), with Pbar stacking (24.9E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7629 at 3:30 AM and put the TeV into shot setup. Operations established store 7630 at 5:10 AM with an AIL of 309.71E30. Operators disabled the NuMI beam at 6:05 AM to allow a cool-off period prior to the scheduled day shift access.

During the day shift, NuMI personnel began their access at 9 AM and ended it by 11:30 AM. RF technicians replaced the modulator for station BRF8. Operators rebooted the MI low-level RF system to resolve a communication issue with RF stations 16, 17, and 18.

During the evening shift, Operations turned off the NuMI power supply V100 to investigate a control card issue. Operations disabled the MTest beam at 6 PM. Booster personnel accessed their ring at 6:20 PM to investigate an airflow problem with BRF14. They came out of access at 7:25 PM.

Wednesday Feb. 24

The midnight shift began with Operations monitoring store 7630 and the stash (386.6E10), with Pbar stacking (21.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7630 at 12:25 AM and put the TeV into shot setup. Operations established store 7631 at 2:25 AM with an AIL of 292.8E30. LRF5 tripped off at 4:25 AM; Linac personnel came in to investigate; repairs lasted the rest of the shift.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster personnel to access and replace a weak RF power amplifier during the morning.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/24/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory